# Neural models of multiscale systems: conceptual limitations, stochastic parametrizations, and a climate application

### v2

@@ IV.2 A reduced order model of surface temperature and radiative fluxes
Given the general problem outlined above, we aim to build a skillful emulator of sea surface temperature and radiative fluxes that allows us to study stationary and perturbed statistics. Building on the previous Sections we consider a functional form of the neural model defined follows:
@@
Where $\mathbf{W}$ is the standard Wiener process. As in the previous Sections, we use a neural network to fit the drift part $(\mathbf{L}+\mathbf{n}(\mathbf{x}))$ and another network to fit the multiplicative noise $\mathbf{\Sigma}(\mathbf{x})$. As before, to improve readability we refer to Appendix A for technical details.
We now face the challenging task of selecting appropriate preprocessing steps to define the coarse-grained dynamics to be modeled. These coarse-graining choices are crucial, as the real climate system involves not only SST and radiative fluxes, but also many other interacting variables, i.e. this is a practical example of the case of a partially observed state vector. We now describe the dataset used for this experiment, along with the preprocessing steps taken. The preprocessing procedure largely follows that described in [104] and we will refer the reader to that work for additional details and clarifications where needed.
IV.2.1 Data, preprocessing and coarse-graining
@@
The focus is on a state-of-the-art coupled climate model GFDL-CM4 [111]. The model offers data from a long control run simulating the coupled climate system for 600 years with constant $\text{CO}_{2}$ fixed at pre-industrial level. The ocean component is the MOM6 ocean model [112] with a horizontal grid spacing of 0.25 and 75 vertical layers. The atmospheric component is the AM4 model [113, 114] with a horizontal grid spacing of roughly 100 km and 33 vertical layers. The model considered in Eq. (10) will be trained on the long control run. The variables considered are two: SST and the global mean net radiative flux at the top of the atmosphere (hereafter, often referred to as TOA). The focus on global mean net radiative flux, rather than its spatiotemporal representation is motivated by the fact that such variable does not show large scale patterns of variability. It is in fact a fast variable in respect to SST and the main signal we are interested in is its global average. The SST field is considered only in the latitudinal range $[-45,45]$, a choice informed by earlier studies highlighting the prominence of radiative feedbacks in the tropics and subtropics [30, 107].
Preprocessing and coarse-graining.
Climate dynamics is not only encompassing the SST and radiative fluxes. Therefore, we aim in building a reduced order model focusing on the slow dynamics and parametrizing unobserved degrees of freedom as noise. We consider the control run and proceed in coarse-graining both the spatial and temporal scales. First, starting from daily data, we compute monthly averages. We then remove the seasonal cycle from both the SST and radiative fluxes, allowing us to focus on seasonal anomalies, rather than periodic signals driven by the annual solar forcing. We then decompose the SST spatiotemporal field in a set of $n$ modes, i.e. empirical orthogonal functions (EOFs). Projecting the computed EOFs onto the original data give us $n$ Principal Components (PCs). In formulae, given the SST field $\mathbf{X}(t)\in\mathbb{R}^{N}$, we decompose it as:
$\begin{split}\mathbf{X}(t)=\sum_{i=1}^{n}x_{i}(t)\mathbf{e}_{i}+\sum_{j=n+1}^{N}y_{j}(t)\mathbf{e}_{j},\end{split}$ (11)
and retain only the first $n=20$ modes $x_{i}(t)$, while the remaining components $y_{j}(t)$ represent the fast modes. $\mathbf{e}$ are orthonormal bases (i.e. the EOFs). Therefore, the time series $x_{i}(t)$ will be used to train the emulator in Eq. (10). These two steps, i.e. one month averaging and the projection over EOFs, attempt to define a new coarse-grained system for the slow variability of the SST field. Note that Penland and Sardeshmukh [4] and Penland (2006) [5] already showed how a similar preprocessing would lead to a skillful linear Markov model of the tropical Pacific SST. These SST principal components are then concatenated with the scalar net radiative flux time series: to simplify the notation, we are going to refer to the state vector of the reduced order model as simply as $\mathbf{x}(t)\in\mathbb{R}^{n+1}$.
Is a 600 year-long control run a long-enough simulation to learn a model of SST and radiative fluxes? Whether this is a sufficiently long simulation depend on whether the time series considered sample enough recurrences, see Cecconi et al. (2012) [31]. In a multiscale system like climate, even in the presence of constant external forcing, we observe processes such as multidecadal oscillations which are only sampled a few times in a 600-year long experiment. Therefore, we make the following choice: we high-pass filter the data with a cut-off frequency of $10^{-1}$ years. This removes all low-frequency variability which we would have little chance to learn from data in the first place, and focus on variability inside 10 years only. This choice involves a trade-off: our model will not be able to capture multidecadal variability and we will restrict ourselves to explore subdecadal variability only. At the same time this leads to an effectively larger sample size than we would have had by keeping all time scales. Additionally, this step also allow us to remove drifts present in the control run of the CM4 run, mainly coming from an ocean component that still did not equilibrate.
@@
In this paper we will focus on testing the ability to “generalize” to unseen climate scenarios: we will fix the SST variable with a precomputed trajectory coming from two, 150 years long forced runs, namely a $1$pctCO2 and 4×4$\times$4 ×CO2, and use this information to reconstruct the change in global mean TOA in these simulations. The $1$pctCO2 is an idealized experiment simulating the climate system under a 1% $\text{CO}_{2}$ increase per year starting from pre-industrial conditions. The 4×4$\times$4 ×CO2 is an idealized experiment simulating the climate system under an abrupt quadrupling of $\text{CO}_{2}$ concentration from its pre-industrial level at time $t=0$. Importantly, considering the $1$pctCO2 and 4×4$\times$4 ×CO2 simulations leads to an additional challenge: the radiative fluxes are directly influenced by an increase in atmospheric $\text{CO}_{2}$. However, our reduced order model only has access to SST and radiative fluxes and it cannot explicitly include $\text{CO}_{2}$ concentration. To address this, we remove the portion of the net TOA radiative flux that is attributable solely to rising $\text{CO}_{2}$. This allow us to isolate and study the component of radiative feedback that is driven by changes in SST. This correction relies on the well established fact that the global-mean radiative forcing from $\text{CO}_{2}$ alone increases approximately logarithmically with its concentration [115, 116]. As is standard, for the 4×4$\times$4 ×CO2 experiment, we remove a standard constant of $8Wm^{-2}$ [117]. For the $1$pctCO2 run we instead apply a time-dependent correction of the form $\alpha\log[C_{t}/C_{0}]$; $C_{0}$ and $C_{t}$ respectively representing the concentration of $\text{CO}_{2}$ in the control run and forced experiments.
We note that considering 1 month averages in TOA fluxes is not arbitrary: the immediate response of TOA fluxes to perturbations in SST is fast (e.g., $\sim 20$ days in tropical large-scale convective overturning motion in the GFDL-CM4 model). Therefore, reconstructing the net radiative fluxes given information of the SST in the two forced runs above is a relatively easy task for this system and it converts to a 1 time step prediction. For this point we refer the reader to the conceptual discussion in Section 7 of [104]. In other words, the preprocessing considered here maximizes the chance of having a data-driven model that correctly reconstruct the TOA variable in unseen climates.
@@ IV.2.2 Stationary statistics and forced responses
Given the $n+1$ dimensional state vector $\mathbf{x}(t)\in\mathbb{R}^{n+1}$ defined in the preprocessing step above, we can now train the emulator defined in Eq. (10) on the long control run of the GFDL-CM4 model. The trained emulator is then used to simulate a 600 years long climate trajectory (same length as the one used in the original model GFDL-CM4). The stationary probability distributions of the first SST mode and of the net radiative flux at the TOA are found to be qualitatively well captured, see Figure 6(a,b). The largest discrepancies between the emulator and the original model are found for the autocorrelation functions. Autocorrelation functions are shown for the first SST mode (largely corresponding to the El Niño Southern Oscillation (ENSO) [118]) and for the TOA variable.
@@
The pattern effect problem discussed in IV.1 concerns identifying a direct causal link between patterns of SST warming and changes in radiative fluxes at the TOA. Before attempting any mechanistic interpretation or physical understanding, we first assess whether the neural model we have built possesses predictive skill in capturing forced responses. Specifically, we test if the neural model can be used to reconstruct the global mean radiative flux from a pre-computed, forced sea surface temperature field in forced experiments. In doing so, we also test an important question in neural models: is the neural network model able to generalize to out-of-distribution predictions? To answer these question we consider the two, 150-years long, forced runs of the GFDL-CM4 climate model introduced in Section IV.2.1. We extract the sea surface temperature from these forced runs and project it onto the set of $n=20$ EOF bases found in the control run. We refer to such forced modes as $\mathbf{x}_{\text{SST}}^{\text{f}}$. We then simulate the neural emulator for 150 years by constraining the SST degrees of freedom to be exactly equal to $\mathbf{x}_{\text{SST}}^{\text{f}}$. Given $\mathbf{x}_{\text{SST}}^{\text{f}}(t)$ the model returns the TOA variable, $x_{\text{TOA}}(t+1)$, at the next time step, starting from initial condition $x_{\text{TOA}}(0)=0$. The results is shown in Figure 6(e,f). In both the case of $1$pctCO2 or 4×4$\times$4 ×CO2 runs, the model is able to reconstruct the change in radiative fluxes at the TOA in a quantitative way. We highlight that our preprocessing allowed us to reduce this reconstruction step into a prediction of the $21$-st entry of the state vector at time $t+1$, given information of the first 20 entries at the previous time $t$. This is only possible because of our focus on 1-month averages: the radiative fluxes at the TOA largely respond already at a sub-monthly time scales to temperature perturbations, therefore allowing us to focus on a direct statistical link $\mathbf{x}_{\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$. We also note that it is perhaps surprising that the neural network is able to respond faithfully even when for nonstationary inputs and with temperature values far out of the training distribution (in physical units, the mean temperature anomaly shifts from a $0K$ to $5K$ from the control in both forced simulations).
@@
Figure 6: Panel (a): stationary distributions obtained for the 1st principal component of the sea surface temperature (SST) field as in the original GFDL-CM4 climate model and as emulated by the neural emulator with multiplicative noise in Eq. (10). Units: nondimensional. Panel (b): same as panel (a) but for the global mean radiative flux at the top of the atmosphere (TOA). Units: $W/m^{2}$. The stationary distributions have been obtained after simulating the neural emulator for 600 years. Second row: same as the first row but for the autocorrelation functions. Third row: reconstruction of the net radiative flux at the TOA by the neural emulator in Eq. (10) after constraining the SST degrees of freedom to follow a precomputed trajectory in two forced experiments in the GFDL-CM4 model. The forced experiments considered are the $1$pctCO2 and 4×4$\times$4 ×CO2, see Section IV.2.1 for details. Note: the net flux at the TOA variable is negative upward (i.e., leaving the planet).
IV.2.3 Interpretability and understanding
@@
In the previous Section we showed that the neural emulator proposed in this study can in fact be used to reconstruct the instantaneous change in radiative feedback at the top of the atmosphere given pre-computed changes in the temperature field. As emphasized before, the task of the emulator in the reconstruction/prediction task reduces to the following mapping: $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\in\mathbb{R}^{n}$, $x_{\text{TOA}}(t)\in\mathbb{R}$ and with initial condition $x_{\text{TOA}}(0)=0$. We now ask the following question: What are the SST patterns that contribute the most to this skillfull prediction? To answer this question, we must go beyond the prediction task presented above and identify which components of the sea surface temperature input contributes the most to changes in radiative flux. We explore this question through the framework of linear response theory, by capturing the direct causal link $x_{j,\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $x_{j,\text{SST}}(t)$ represents the $j$-th component of the surface temperature field at time $t$.
We briefly describe the strategy to infer this direct causal link with response theory, and refers to Section II.2 for details. We consider the neural emulator in Eq. (10) and iteratively perturb the $j$-th sea surface temperature component of the state vector at time $t=0$ with a small impulse perturbation $x_{j}(0)+\Delta_{j}\delta(t)$. $\delta(t)$ being a Dirac delta and $\Delta_{j}$ being a perturbation of very small amplitude. The amplitude $\Delta_{j}$ depends on the component $x_{j,\text{SST}}$ itself (Section II.2). We then compute the ensemble mean response of the net radiative flux at the TOA at the next time step $t=1$, referred here as $\delta\langle x_{\text{TOA}}(1)\rangle$. To estimate the ensemble mean response we consider $10^{3}$ initial conditions. We write this response as follows:
$\begin{split}R_{\text{TOA},j}(1)=\frac{\delta\langle x_{\text{TOA}}(1)\rangle}{\Delta_{j}\delta(t)},\end{split}$ (12)
ad remind the reader that the index $j$ only considers the SST degrees of freedom. Therefore, the response at 1-time-step $R_{\text{TOA},j}(1)$ defines the direct causal link $x_{j,\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$ between the $j$-th degree of freedom in SST and the net radiative flux at the TOA.
The vector $\mathbf{R_{\text{TOA}}}(1)\in\mathbb{R}^{n}$ is projected back to the higher, original $N$-dimensional space. To do so we consider the previously computed EOFs, i.e. the orthonormal vectors $\mathbf{e}\in\mathbb{R}^{n,N}$ and project the 1-time-step responses as follows:
$\begin{split}\mathbf{S}=\mathbf{R_{\text{TOA}}}(1)\cdot\mathbf{e}=\sum_{j}^{n}R_{\text{TOA},j}(1)~{}\mathbf{e}_{j},\end{split}$ (13)
@@
where $\mathbf{S}\in\mathbb{R}^{N}$ is a sensitivity map defined in the high-dimensional $N$ dimensional space. Such map approximates at each grid point $i$ in the ocean, the sensitivity $\frac{\partial x_{\text{TOA}}}{\partial x_{\text{SST},i}}$, where $x_{\text{TOA}}$ is the global mean radiative flux at the TOA and $x_{\text{SST},i}$ is the SST at grid point $i$.
The sensitivity map $\mathbf{S}$ obtained using the neural emulator in Eq. (10) is shown in Figure 7. The obtained sensitivity map is in qualitative agreement with previous results in response theory (Figure 2(a) in [104]), with other data-driven approaches (see for example Figure 7(b) in [119] and Figure 2 in [120]), but, most importantly, with the sensitivity map obtained by directly perturbing the SST boundary condition of atmosphere-only climate models (see third row of Figure 2 in [29], see Figure 5(a) in [30]). Note that in the latter case, the maps exhibit fewer spatial features due to the coarse resolution of the perturbation patches, which span tens of degrees. The marked feature of this sensitivity map is the sensitivity dipole found the tropical Pacific, with negative values on the western side of the basin and positive values on the eastern side, see for example [30]. This means that an initial warming in the eastern Pacific basin will result in a direct, positive radiative feedback, further amplifying the imposed temperature change; the opposite is true for the western Pacific. The mechanistic explanations for such mechanisms can be found for example in [121, 110] and we refer the interested reader to these papers for details. In brief, the positive sensitivity in $\mathbf{S}$ in the eastern Pacific is associated with a high-climatology of low-level clouds: a warming perturbation is expected to decrease the cloud cover, leading to anomalies in incoming shortwave radiation and therefore a positive sensitivity. On the other hand, the western Pacific is a region dominated by deep convective phenomena. Because of this, a surface temperature warming in the western Pacific basin tend to travel vertically through the troposphere and then horizontally through gravity waves. This process strengthen the inversion layer in the eastern Pacific basin (and in other regions with high-climatology of low-level clouds), leading to a larger low-level cloud fraction and increasing the reflected shortwave at the top of the atmosphere.
To conclude this Section, the analysis in Figure 7 confirms that the good reconstruction observed in the forced regime (Figure 6(e,f)) arises for the physically correct reasons.
@@
Figure 7: Sensitivity map obtained using the neural emulator in Eq. (10). At each grid point $i$ we are plotting the 1 time step response in the global mean, net radiative flux at the top of the atmosphere (TOA) driven by an impulse perturbation in the sea surface temperature (SST) at that grid point. Note: the computation is done in a 21 dimensional space spanned by a set of empirical orthogonal functions and this map is a projection of the results obtained in the low-dimensional space. The response has been obtained by averaging over an ensemble of $10^{3}$ initial conditions.

### v3

@@ IV.2 A reduced order model of surface temperature and radiative fluxes
We aim to build a skillful emulator of large-scale sea surface temperature (SST) and radiative fluxes capable of capturing both stationary and perturbed statistics. Building on the previous sections, we adopt the following functional form for the neural model:
@@
Where $\mathbf{W}$ is the standard Wiener process. As before, to improve readability we refer to Appendix C for technical details.
As outlined throughout the paper, a central challenge lies in selecting preprocessing steps to define the coarse-grained dynamics $\mathbf{x}$ to be modeled. These choices are critical for identifying the proper variables, since the climate system involves not only SST and radiative fluxes but also many other interacting variables. This is thus a practical instance of a partially observed state vector. Below, we describe the dataset and preprocessing procedure, which largely follows [102], with further details provided in that work.
IV.2.1 Data, preprocessing and the choice of proper variables
@@
We consider outputs from the state-of-the-art coupled climate model GFDL-CM4 [103], which provides a 600-year pre-industrial control run with constant $\text{CO}_{2}$. The ocean component is MOM6 [104] with 0.25 horizontal resolution and 75 vertical layers, and the atmosphere is AM4 [105, 106] with  100 km resolution and 33 vertical layers. The model in Eq. (11) is trained on this control run. We consider two climate fields: sea surface temperature (SST) and the global mean net radiative flux at the top of the atmosphere (TOA). The SST field is restricted to latitudes $45^{\circ},45^{\circ}]$, following earlier studies that highlight the importance of radiative feedbacks in the tropics and subtropics [39, 97].
Choosing the proper variables.
Climate dynamics involves many interacting fields beyond SST and radiative fluxes. Our goal is a reduced-order model that captures the coarse-grained dynamics while representing unresolved processes as noise. We coarse-grain the GFDL-CM4 control run in space and time, averaging daily data to monthly means and removing the seasonal cycle to focus on anomalies rather than periodic signals driven by solar forcing. Considering 1 month averages in TOA fluxes is not arbitrary: the immediate response of TOA fluxes to perturbations in SST is fast (e.g., $\sim 20$ days in tropical large-scale convective overturning motion in the GFDL-CM4 model) [102]. This averaging filters out higher-frequency processes that mediate surface–TOA coupling, yielding a system that is more Markovian. Given so, direct causal links between TOA fluxes at SST patterns can be detected as one time step responses; see Appendix A. We then high-pass filter with a cut-off of $10^{-1}$ years, removing (i) low-frequency (multidecadal) oscillations, which are poorly sampled even in 600-year runs, and (ii) slow drifts from incomplete ocean equilibration. This choice sacrifices the ability to capture multidecadal processes but ensures robust sampling of subdecadal statistics, yielding a larger effective sample size. Implicitly, the choice assumes that the response of the net radiative flux to SST perturbations occurs within a 10-year window. Next, we decompose the SST field $\mathbf{X}(t)\in\mathbb{R}^{N}$ into empirical orthogonal functions (EOFs):
$\mathbf{X}(t)=\sum_{i=1}^{n}x_{i}(t)\mathbf{e}_{i}+\sum_{j=n+1}^{N}y_{j}(t)\mathbf{e}_{j},$ (12)
and retain the first $n=20$ modes $x_{i}(t)$. The combined steps of monthly averaging, high-pass filtering and projection over EOF bases define a coarse-grained system for the slow SST variability. This choice follows earlier work showing the skill of linear Markov models built on similar preprocessing [4, 5]. For radiative fluxes, we focus on the global mean TOA net flux, as the TOA flux exhibits little spatial structure and we are mainly interest in its average value. The state vector is then defined as $\mathbf{x}(t)\in\mathbb{R}^{n+1}$, consisting of the $n$ SST principal components concatenated with the global mean TOA flux.
Forced simulations
@@
We test the model’s ability to reconstruct changes in global-mean TOA radiative flux given prescribed SST trajectories from two 150-years long forced runs: 11pctCO2 and $4\timesCO$2. The 11pctCO2 experiment imposes a 1% $\text{CO}_{2}$ increase per year starting from pre-industrial conditions, while the $4\timesCO$2 experiment abruptly quadruples $\text{CO}_{2}$ at $t=0$. A complication is that TOA fluxes respond directly to rising $\text{CO}_{2}$, whereas our reduced-order model only uses SST and fluxes and does not explicitly include $\text{CO}_{2}$. To address this, we remove the portion of change of the net TOA radiative flux that is attributable solely to rising $\text{CO}_{2}$. This allow us to isolate and study the component of radiative feedback that is driven by changes in SST. This correction exploits the logarithmic dependence of global-mean forcing on $\text{CO}_{2}$ concentration [107, 108]. As is standard, we subtract $8\text{Wm}^{-2}$ in the $4\timesCO$2 run [109], and apply a time-dependent correction $\alpha\log(C_{t}/C_{0})$ in the 11pctCO2 run, where $C_{t}$ and $C_{0}$ denote $\text{CO}_{2}$ concentrations in the forced and control simulations.
@@ IV.2.2 Stationary statistics and forced responses
Given the $n+1$ dimensional state vector $\mathbf{x}(t)\in\mathbb{R}^{n+1}$ defined above, we train the emulator in Eq. (11) on the long control run of GFDL-CM4. The trained model is then used to simulate a trajectory of equal length (i.e. 600 years at monthly time step). The stationary probability distributions of the leading SST mode and the net radiative TOA flux are well reproduced (Fig. 6a,b). The main discrepancies arise in the autocorrelation functions, shown for the first SST mode (closely linked to ENSO [110]) and the TOA flux.
@@
The pattern-effect problem (Section IV.1) concerns identifying causal links between SST warming patterns and TOA flux changes. Before using our emulator for Green’s function–like studies to investigate the pattern-effect, we first assess its skill as a reconstruction tool. Specifically, we reconstruct global-mean TOA flux from prescribed SST trajectories in the two forced experiments, i.e. 11pctCO22 and $4\timesCO$22. Forced SSTs are projected onto the first $n=20$ EOFs from the control run, yielding coarse-grained trajectories $\mathbf{x}_{\text{SST}}^{f}(t)$, $f$ denotes “forced.” Given $\mathbf{x}_{\text{SST}}^{f}(t)$, the emulator returns the TOA variable $x_{\text{TOA}}(t+1)$ at the next time step, initialized with $x_{\text{TOA}}(0)=0$. Results (Fig. 6e,f) show that the emulator largely reproduces TOA flux changes in both scenarios. This reconstruction is possible because we focus on one-month averages, which reduce the time-dependent TOA response, driven by an atmospheric reorganization at sub-monthly scales, into a one-step prediction. For the 1pctCO2 experiment, the linear trend in net TOA flux is $0.043Wm^{-2}yr^{-1}$ in GFDL-CM4 and $0.034Wm^{-2}yr^{-1}$ in the emulator. We then compute correlations between reconstructed and simulated radiative flux at the TOA, after removing a linear trend in 1pctCO2 and an initial 50-year transient in 4xCO2. At monthly resolution, correlations are $r=0.37$ (1pctCO2) and $r=0.43$ (4xCO2). If yearly averages are considered (as in the current literature, see e.g. [38]) these increase to $r=0.70$ (1pctCO2) and $r=0.74$ (4xCO2). All correlations are significant at the $5$ sigma level under a null hypothesis of zero correlation, tested via bootstrap resampling with replacement.
@@
Figure 6: Panel (a): stationary distributions obtained for the 1st principal component of the sea surface temperature (SST) field as in the original GFDL-CM4 climate model and as emulated by the neural emulator with multiplicative noise in Eq. (11). Units: nondimensional. Panel (b): same as panel (a) but for the global mean radiative flux at the top of the atmosphere (TOA). Units: $W/m^{2}$. The stationary distributions have been obtained after simulating the neural emulator for 600 years. Second row: same as the first row but for the autocorrelation functions. Third row: reconstruction of the net radiative flux at the TOA by the neural emulator in Eq. (11) after constraining the SST degrees of freedom to follow a precomputed trajectory in two forced experiments in the GFDL-CM4 model. The forced experiments considered are the 11pctCO2 and $4\timesCO$2, see Section IV.2.1 for details. Note: the net flux at the TOA variable is negative upward (i.e., leaving the planet).
IV.2.3 Interpretability and understanding: ensemble mean and variance responses
@@
In the previous section, we showed that the reduced-order neural emulator can reconstruct the net radiative flux at the top of the atmosphere (TOA) from prescribed changes in the sea surface temperature (SST) field. This amounts to the mapping $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $\mathbf{x}_{\text{SST}}(t)\in\mathbb{R}^{n}$, $x_{\text{TOA}}(t)\in\mathbb{R}$, and $x_{\text{TOA}}(0)=0$. We now ask the following question: What are the SST patterns that contribute the most to this skillfull prediction? Addressing this question requires moving beyond prediction to identify the causal influence of individual SST components on radiative flux. We frame the problem in terms of linear response theory, focusing on the direct link $x_{j,\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $x_{j,\text{SST}}(t)$ denotes the $j$-th SST component.
Response in ensemble mean.
To quantify this link, we perturb the emulator in Eq. (11) following the procedure in Appendix A. Briefly, at $t=0$, we apply a small impulse $\Delta_{j}=10^{-2}\sigma_{j}$ to the $j$-th SST component, with $\sigma_{j}$ its stationary standard deviation. We then compute the ensemble-mean response of the net TOA flux at $t=1$, denoted $\delta\langle x_{\text{TOA}}(1)\rangle$. This quantity, referred to as $R^{\mu}_{\text{TOA},j}(1)$, represents the ensemble-mean response ($\mu$ indicates the mean) and is estimated over $10^{6}$ initial conditions to ensure robust statistics
Response in ensemble variance.
While the ensemble-mean response is informative, it captures only part of the adjustment. Large-scale climate responses can involve shifts of probability distributions. To partially account for this, we quantify the response in the variance of the net TOA radiative flux to surface warming perturbations. As above, the estimate is based on an ensemble of $10^{6}$ trajectories. For consistency in physical units, we express the response in terms of standard deviation rather than variance, denoting it by $R^{\sigma}_{\text{TOA},j}(1)$ ($\sigma$ indicates the standard deviation). This quantity measures the change in ensemble spread following a perturbation of the $j$-th SST component.
Conceptually, this procedure provides a computationally efficient analogue of atmospheric-only Green’s function experiments, while shifting the focus from individual trajectories to full probability distributions.
Sensitivity maps.
The vectors of one-time-step responses in mean and standard deviation, $\mathbf{R}^{\mu}$ and $\mathbf{R}^{\sigma}$, are projected back to the full $N$-dimensional SST field using the EOFs, $\mathbf{e}\in\mathbb{R}^{n,N}$:
$\begin{split}\mathbf{S}^{\mu,\sigma}=\mathbf{R^{\mu,\sigma}_{\text{TOA}}}(1)\cdot\mathbf{e}=\sum_{j}^{n}R^{\mu,\sigma}_{\text{TOA},j}(1)~\mathbf{e}_{j},\end{split}$ (13)
@@
where $\mathbf{S}^{\mu},\mathbf{S}^{\sigma}\in\mathbb{R}^{N}$ represent sensitivity maps in mean and standard deviation, respectively. At each ocean grid point $i$, these maps approximate the sensitivities $\frac{\partial x^{\mu,\sigma}_{\text{TOA}}}{\partial x_{\text{SST},i}}$, where $x^{\mu}_{\text{TOA}}$ and $x^{\sigma}_{\text{TOA}}$ respectively represent the ensemble mean and standard deviation of the mean, net radiative flux and $x_{\text{SST},i}$ is the SST at location $i$.
The sensitivity maps in mean ($\mathbf{S}^{\mu}$) and standard deviation ($\mathbf{S}^{\sigma}$) are shown in Figure 7. The obtained sensitivity map $\mathbf{S}^{\mu}$ is in qualitative agreement with previous results in response theory (Figure 2(a) in [102]), with other data-driven approaches (see Figure 7(b) in [111] and Figure 2 in [112]), but, most importantly, with the sensitivity map obtained by directly perturbing the SST boundary condition of atmosphere-only climate models (see third row of Figure 2 in [38], see Figure 5(a) in [39]). In the latter case, the maps exhibit fewer spatial features due to the coarse resolution of the perturbation patches (spanning tens of degrees) in numerical models experiments.
@@
Figure 7: Sensitivity maps obtained using the neural emulator in Eq. (11). At each SST grid point $i$, we show the one–time-step response in the ensemble mean ($\mathbf{S}^{\mu}$) and ensemble standard deviation ($\mathbf{S}^{\sigma}$) of the global-mean net radiative flux at the TOA following an impulse perturbation in SST at that location. The computation is carried out in a 21-dimensional reduced space spanned by empirical orthogonal functions, and the maps represent the projection of these low-dimensional results back to the full grid. Responses are estimated from an ensemble of $10^{6}$ initial conditions.
The dominant feature of $\mathbf{S}^{\mu}$ is a dipole in the tropical Pacific: negative sensitivity in the west and positive sensitivity in the east [39]. This implies that eastern Pacific warming produces a positive radiative feedback, while western Pacific warming produces the opposite. The mechanisms are well established [113, 100]: in the east, high climatological low-level cloud cover decreases under warming, reducing reflected shortwave radiation and yielding positive sensitivity. In the west, deep convection transports surface warming vertically and then horizontally via gravity waves, strengthening the inversion over the eastern Pacific and other stratocumulus regions, increasing low-level cloud fraction and reflected shortwave radiation, thus producing a negative feedback. This analysis confirms that the accurate reconstructions in the forced regime (Fig. 6e,f) arise from physically consistent mechanisms, and that coarse-grained neural models can recover such causal linkages.
Finally, the variance response $\mathbf{S}^{\sigma}$ (Fig. 7b) highlights regions where perturbations alter the spread of radiative responses across ensemble members. Perturbations in stratocumulus-dominated regions (northeast Pacific, off the coast of South America) yield negative variance responses: the ensemble distribution of net radiative flux narrows relative to the control, highlighting small uncertainties in the mean response studied above. By contrast, surface warming perturbations in the central and western Pacific lead to positive variance responses, reflecting a broader spread of the net radiative flux distribution after SST perturbations. To our knowledge, such variance responses have not been quantified before. Unlike the response in the mean, which has been examined extensively and shows broad agreement across atmospheric models [38], the response in ensemble variance has no established benchmark. Because it reflects changes in variability, it may differ significantly across atmospheric models. Moreover, estimating such responses requires very large ensembles, which are prohibitively expensive in conventional climate models. For this reason, the results presented here should be viewed as an exploratory step: they highlight potentially important regional structures, but a full climate interpretation will require systematic intercomparison and validation against other modeling frameworks, such as novel data-driven tools to estimate response through the fluctuation-dissipation theorem [114].